OpenStudy (anonymous):

how do you factor an equation when the poduct need to be negative but when you make one number negative it no longer is the sum for the other part of the equation. for exanple... 14x^2+3x-2=0

OpenStudy (anonymous):

here factor 14 = 7 * 2 -2 = -2 * 1 or -1 * 2 we need + 3 so its 7* 1 + 2 * -2 = +3 so the -2 and + 1 is required so factors are (7x -2 )(2x + 1)

OpenStudy (anonymous):

good job jimmyrep

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

its a matter of trial and error 14 also equals 14 * 1 - but this would not combine with factors of -2 to get +3. it can be a bit tricky but becomes a lot easier with practice.

OpenStudy (anonymous):

hope this has helped

OpenStudy (anonymous):

when factoring, does it matter which numbers you chose and which to be negative or positive as long as they equal the sum of the "x"'s variable and the product of the number??

OpenStudy (anonymous):

can you put spaces or commas btw the numbers bc the 14=7*2-2 is not 14.